for the opening day of a carnival, 800 admission tickets were sold. the receipts totaled $3775. tickets for children cost $3 each, tickets for adults cost $8 each, and tickets for senior citizens cost $5 each. there were twice as many childre's tickets sold as adult tickets. How many of each type of ticket were sold?
Answer by lwsshak3(11628)

for the opening day of a carnival, 800 admission tickets were sold. the receipts totaled $3775. tickets for children cost $3 each, tickets for adults cost $8 each, and tickets for senior citizens cost $5 each. there were twice as many childre's tickets sold as adult tickets. How many of each type of ticket were sold?
**
let x=number of adult tickets sold
2x=number of children tickets sold
800-x-2x
800-3x=number of senior tickets sold
..
8x+3(2x)+5(800-3x)=3775
8x+6x+4000-15x=3775
-x=-225
x=225
2x=450
800-3x=800-675=125
...
ans:
Tickets sold:
Adults= 225
Children: 450
Seniors: 125